William A. Stein Curriculum Vitae

(617)308-0144  ·  was@math.harvard.edu  ·  http://modular.fas.harvard.edu

Employment · Harvard University, Benjamin Peirce Assistant Professor, 2001–present. · NSF Postdoctoral Fellowship, 2000–2001, 2003–2004. · Defense consulting (see http://dssg.ida.org), 2002–2003.

Education · University of California at Berkeley, Ph.D., May 2000, Explicit approaches to modular abelian varieties.

Publications 17. Constructing Elements in Shafarevich-Tate Groups of Modular Motives (19 pages), with N. Dummigan and M. Watkins, to appear in “Number theory and algebraic geometry—to Peter Swinnerton-Dyer on his 75th birthday”, Ed. M. Reid and A. Skorobogatov.

16. Approximation of Infinite-Slope Modular Eigenforms By Finite-Slope Eigenforms (13 pages), with R. Coleman, to appear in the Dwork Proceedings.

15. J_{1}(p) has connected fibers, with B. Conrad and B. Edifxhoven (70 pages), to appear in Documenta Mathematica.

14. Shafarevich-Tate Groups of Nonsquare Order (12 pages), to appear in the 2002 MCAV proceedings.

13. Visible Evidence for the Birch and Swinnerton-Dyer Conjecture for Rank 0 Modular Abelian Varieties (30 pages), with A. Agashe, to appear in Mathematics of Computation.

12. A Database of Elliptic Curves–First Report (10 pages), in ANTS V proceedings, Sydney, Australia, 2002.

11. Visibility of Shafarevich-Tate Groups of Abelian Varieties (19 pages), with A. Agashe, J. Number Theory, 97 (2002), no. 1, 171–185.

10. Transportable modular symbols (12 pages), with H. Verrill, LMS J. Comput. Math., 4 (2001), 170–181.

9. Appendix to Lario and Schoof’s Some computations with Hecke rings and deformation rings (1 page), with A. Agashe, to appear in Experiment. Math.

8. There are genus one curves over Q of every odd index (9 pages), J. Reine Angew. Math. (Crelle’s Journal).

7. Component groups of optimal quotients of Jacobians (20 pages), with B. Conrad, , Math. Res. Lett., 8 (2001), no. 5-6, 745–766.

6. The field generated by the points of small prime order on an elliptic curve (7 pages), with L. Merel, Internat. Math. Res. Notices, 2001, no. 20, 1075–1082.

5. An introduction to computing modular forms using modular symbols (10 pages), to appear in an MSRI proceedings volume.

(617)308-0144  ·  was@math.harvard.edu  ·  http://modular.fas.harvard.edu

4. Mod 5 approaches to modularity of icosahedral Galois representations (18 pages), with K. Buzzard, Pacific J. Math, 203 (2002), no. 2, 265– 282.

3. Lectures on Serre’s conjectures (95 pages), with K. A. Ribet, in Arithmetic Algebraic Geometry, IAS/Park City Math. Inst. Series, Vol. 9, 143–232.

2. Component groups of quotients of J_{0}(N) (9 pages), with D. Kohel, Proceedings of the 4th International Symposium (ANTS-IV), Leiden, Netherlands, July 2–7, 2000, Springer, 2000.

1. Empirical evidence for the Birch and Swinnerton-Dyer conjectures for modular Jacobians of genus 2 curves (22 pages), with E. V. Flynn, F. Lep-révost, E. F. Schaefer, M. Stoll, J. L. Wetherell, Math. of Comp. 70 (2001), no. 236, 1675–1697.

Books in progress · Elementary Number Theory and Elliptic Curves (250 page book), under contract for Springer-Verlag’s UTM series. · Lectures on Modular Forms and Galois Representations (200 page book), with K. A. Ribet, intended for Springer-Verlag’s Universitext series.

Computation · Author of the modular forms and modular symbols parts of Magma (Three visits to Sydney, Australia to work with the Magma group.) · The Modular Forms Database:

http://modular.fas.harvard.edu/Tables/.

Equipment Grants Each of the following grants provided substantial computing resources to many members of the department: · Sun Education Grant ($70K Sun Fire V480 server), 2003. · 12 Processor Cluster ($20K from W. R. Hearst III and Harvard), 2002. · Vice Chancellor Research Grant (six-processor cluster), 1999.

Fellowships · NSF Postdoctoral Research Fellowship, 2000–2001, 2003-2004. · Clay Mathematics Institute Liftoff Fellowship, Summer 2000. · Two years support from competitive fellowships at Berkeley.

Teaching Harvard University · Modular Abelian Varieties, graduate course, Fall 2003. · Freshman Seminar on Elliptic Curves, Spring 2003. · Elementary Number Theory, Fall 2001 and 2002. · Linear Algebra, Fall 2001 and Spring 2002.

Recent Seminar organization

(617)308-0144  ·  was@math.harvard.edu  ·  http://modular.fas.harvard.edu

· Harvard Modular Curves Seminar, 2000–2003.

NSF Sponsored IAS/Park City Mathematics Institute · Teaching Assistant, Summer 1999. Led problem sessions and prepared notes for Ken Ribet’s course on Serre’s conjectures for advanced number theory graduate students.

University of California at Berkeley

· Curriculum Development, Fall 1997–Summer 1998. Developed curriculum materials and Matlab software for workshop-based calculus and linear algebra courses at UC Berkeley. · Instructor, Summer 1997. Discrete Mathematics. Evaluation average: 6.1 out of 7. · Teaching Assistant, Fall 1995–Spring 1997.

Northern Arizona University

· Teaching, 1994–1995. Taught complete undergraduate courses on elementary combinatorics and algebra.

Personal 345 Harvard Street, #4D Cambridge, MA 02138 US Citizen, born in Santa Barbara, California on February 21, 1974. Phone: (617)308-0144

References

Professor Robert F. Coleman Department of Mathematics #3840 University of California Berkeley, CA 94720-3840

(510) 642-5101 coleman@math.berkeley.edu Professor Benedict Gross Department of Mathematics Harvard University Cambridge, MA 02138

(617) 495-2172

gross@math.harvard.edu

Professor Barry C. Mazur Department of Mathematics Harvard University Cambridge, MA 02138

(617) 495-2171 ext. 512 mazur@abel.math.harvard.edu Professor Kenneth A. Ribet Department of Mathematics #3840 University of California Berkeley, CA 94720-3840

(510) 642-0648

ribet@math.berkeley.edu